# Defining Neural Network Architecture through Polytope Structures of Dataset

### v1

@@ 3.1 Data geometry-dependent bounds on widths
for some constant $M$. Then $\mathcal{N}({\bm{x}})=1$ for all ${\bm{x}}\in C$, and we can prove that $d\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}0pt\rightarrow 1$ is a feasible architecture on the polytope, by adjusting the value of $M$. The rest of the proof can be found in Appendix E.1. ∎
@@
Theorem 3.4 can be further used to derive upper bounds on width of feasible architecture in terms of topological quantities of the given dataset. First, we investigate such bounds and feasible architecture on a simplicial complex, which is widely used in topological data analysis (TDA) to investigate the topological structure of a given point clouds.
@@
The proof can be found in Appendix E.3. Theorem 3.5 reveals that the width $d_{1}$ is bounded by in terms of the number of facets $k$ of the provided simplicial complex. From a geometric perspective, it is generally intuitive that a smaller number of facets suggests a simpler structure of the simplicial complex. This notion is mathematically expressed in (3.5), which suggests that the first value in (3.5) results in $d_{1}\lesssim\frac{k}{2}(d+3)$, which magnifies as $k$ increases. Similarly, when the maximal dimension $J$ is smaller than $\frac{d}{2}$ and $k$ is fixed, the summation in the second term in (3.5) reduces to $d_{1}\lesssim(d+1)\left(k\frac{J+2}{d-J}+2\right)$, which rapidly diminishes as $J$ decreases. This analysis demonstrates that a smaller dimension $J$ demands smaller widths, which aligns with the intuition that a low-dimensional manifold could be approximated with a small number of neurons.
The result in Theorem 3.4 can be further leveraged to ascertain a neural network architecture with width bounds defined in terms of the Betti numbers. Recall that Theorem 3.4 offers an upper bound on widths when $\mathcal{X}$ can be depicted as a difference between groups of convex sets. Expanding on this, assuming the polytope-basis cover consists of prismatic polytopes, we can derive a bound for network architecture in relation to its Betti numbers. This concept is further explained in the following theorem.
@@
Let $\mathcal{X}=\mathcal{X}_{+}\cup\mathcal{X}_{-}$ be a union of two disjoint topological spaces, where $\mathcal{X}_{-}$ can be separated from $\mathcal{X}_{+}$ by disjoint bounded prismatic polytope having at most $0pt$ faces. Let $\beta_{k}$ be the $k$-th Betti number of the polytope-basis cover. Then, the following three-layer architecture
@@
The proof is in Appendix E.4. One of the important implications of Theorem 3.6 is the upper and lower bounds on network widths in terms of the Betti numbers of $\mathcal{X}$, which reveals the interplay between the topological characteristics of the dataset and network architectures. In Appendix, we also show in Proposition F.2 that topological property alone cannot determine the feasible architecture, highlighting the significance of prismatic polytopes assumption in Theorem 3.6. In other words, the result in Proposition F.2 implies that the geometrical assumptions in this theorem are indispensable to connecting topological features with bounds on the network widths. Furthermore, the lower bound on widths (4) shows that the sum of product of widths should be greater than the sum of Betti numbers. This finding also confirms the increased significance of widths in deeper layers as compared to earlier ones, highlighting the advantageous impact of depth in network architecture.
@@ 3.2 Polytope-basis cover search algorithm
Figure 4: A polytope-basis cover derived from a trained three-layer ReLU network defined in (6), obtained from Algorithm 1. The decision boundary and activation boundaries333 For the formal definition of decision boundary (DB) and activation boundaries (ABs), see Definition C.1 in Appendix C.of the trained network are depicted in (a). Each polytope corresponding to $a_{j}=+1$ and $a_{j}=-1$ is illustrated in (b) and (c), respectively. The result constitutes a polytope-basis cover of the swiss roll dataset.
@@
$\mathcal{T}_{j}({\bm{x}})$ $:=1+\sum_{k=1}^{0pt_{j}}v_{jk}\sigma({\bm{w}}_{jk}^{\top}{\bm{x}}+b_{jk}),\quad\forall v_{jk}<0,$ (5)
$\mathcal{N}({\bm{x}})$ $:=-\frac{1}{2}+\sum_{j=1}^{J}a_{j}\sigma(\mathcal{T}_{j}({\bm{x}})),\quad\forall a_{j}\in\{\pm 1\}.$ (6)
For a given dataset ${{\mathcal{D}}}=\{({\bm{x}}_{i},y_{i})\}$, suppose $\mathcal{N}$ satisfies
$\sigma(\mathcal{T}_{j}({\bm{x}}_{i}))=0\;\textrm{ or }\;1,\quad\forall{\bm{x}}_{i}\in{{\mathcal{D}}},\;\forall j\in[J].$ (7)
Then, the collection of polytopes $\{C_{j}\}_{j\in[J]}$, defined by $C_{j}:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}_{j}({\bm{x}})=1\}$, becomes a polytope-basis cover of ${{\mathcal{D}}}$ whose accuracy is same with $\mathcal{N}$.
Algorithm 1 Compressing algorithm.
@@
The proof of this theorem can be found in Appendix E.5. This theorem establishes that if a trained three-layer network $\mathcal{N}$, defined in (6), satisfies the condition outlined in (7), then we can derive a polytope-basis cover of the training dataset ${{\mathcal{D}}}$ from $\mathcal{N}$. Below, we outline two strategies we employed to satisfy both (5) and (6) conditions.
First, to meet (5), we initialize the network weights to satisfy
@@
Then, according to the implicit bias of gradient descent established by Du et al. (2018, Theorem 2.1), as stated in Proposition F.5, it is ensured that $v_{jk}<0$ for all $j\in[J]$ and $k\in[m]$ on the gradient flow. This satisfies the first condition.
Second, to achieve (7), we introduce a novel approach called the “compressing algorithm.” This algorithm aims to ‘compress’ a given two-layer network $\mathcal{T}_{j}$ defined in (5) to identify a minimal convex polytope representation. The process is precisely outlined in Algorithm 1.
More precisely, the algorithm operates in two main steps: 1) identifying and removing a neuron that do not significantly influence the decision boundary, and 2) amplifying the weights to delineate the faces of the decision boundary polytope. We illustrate the functionality of the algorithm in Figure 3. In Figure 3(a), the neuron indicated by the red arrow is deemed non-essential to the decision boundary and is consequently eliminated in (b). Subsequently, by scaling the weights $(v_{k},{\bm{w}}_{k},b_{k})$ by a factor $\lambda_{scale}>1$, the decision boundary shrinks into a convex polytope with a number of faces equal to the width of $\mathcal{T}$ (depicted in Figure 3(c)).
Figure 5: Experimental verification of convergence of gradient descent. Two columns exhibit the shape of two topological spaces, which are ‘two triangles’ and ‘a hexagon with a pentagon hole’. The second and third row show the converged networks by gradient descent under the BCE loss, and the MSE loss, respectively. The first layer (blue) and second layer (red) represent the hyperplane and polytopes, respectively, as discussed in Section 3.
Note that a single execution of Algorithm 1 may not immediately yield the network satisfying (7). However, we prove in Proposition C.2 that by iterating this algorithm a sufficient number of times, the output of the algorithm always satisfies both (5) and (7), making it suitable for the application of Theorem 3.7. Therefore, we apply the algorithm once every thousand iterations during the gradient descent optimization process. Further details on the complete training process can be found in Appendix C.1.
Consequently, we present a polytope-basis cover of the swiss roll dataset derived from a converged three-layer network in Figure 2. The polytopes comprising the resulting cover are visualized in Figure 2 (b) and (c). This outcome demonstrates the polytope-basis cover inherent in the trained network.
Lastly, we mention that we further propose two alternative methods for obtaining a polytope-basis cover in Appendix C. Specifically, one approach involves deriving such a cover from a trained two-layer network (Algorithm 3), while the other method entails training several neural networks (Algorithm 4). Further details and results of these additional algorithms are provided in Appendix C.
@@ 4 Experimental Results
In Section 3, we studied the relationship between the dataset geometry and neural network architectures. In this section, we provide two empirical results: 1) gradient descent indeed converges to the networks we unveil, and 2) we can investigate the geometric features of real-world datasets through our proposed algorithm.
@@ 4.1 Convergence on polytope-basis covers
For the first dataset $\mathcal{X}_{1}$, Theorem 3.5 suggests that $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}6\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}2\rightarrow 1$ is a feasible architecture on $\mathcal{X}_{1}$. To facilitate a clearer examination of weight vectors in each layer, we plot the activation boundaries3 in blue (the 1st hidden layer) and red (the 2nd hidden layer) colors, where the gray color denotes the decision boundary of the converged network. Moreover, the weight vectors in the first layer accurately enclose the two triangles, reflecting the geometrical shape of $\mathcal{X}_{1}$. Similarly, for the second dataset $\mathcal{X}_{2}$, Theorem 3.4 suggests that $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}11\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}2\rightarrow 1$ is a feasible architecture . Specifically, the eleven neurons in the first layer correspond to the eleven hyperplanes delineating the boundaries of the outer hexagon and the inner pentagon, while two neurons in the second hidden layer align with the two polygons. These outcomes precisely correspond to our network constructions depicted in Figure 2(b, c).
We conclude this section by providing theoretical insights into the convergence behavior of gradient descent. In Appendix D, utilizing our explicit construction of neural networks, we demonstrate that gradient descent reliably converges to global minima when the decision boundary of the network is initialized close to the target polytopes (see Theorem D.3). The specific conditions governing the initialization region are described in terms of the geometric distribution of the dataset. For a thorough understanding and precise statements, we refer readers to Appendix D.
@@ 4.2 Polytope-basis cover for real datasets
Figure 6: Illustration of a polytope-basis cover of the real datasets. (a) The class $\{0\}$ of MNIST can be separated by a single convex polytope with four faces, while the complement class $\{0\}^{c}$ can be separated with three faces. (b) The class $\{0\}$ of Fashion-MNIST can be separated by the difference of two polytopes, one of which contains similar images.
@@
Our empirical results are summarized in Table 1. Each column in the table corresponds to a class in the dataset, where each row presents the type of the class. The values in the table denotes the number of polytopes and their faces (we use notation $a+b$ to denote two polytopes with $a$ and $b$ faces, respectively). For instance, the value in the first row and the first column shows that the $\{0\}$ class images in MNIST dataset can be distinguished from other classes by a single convex polytope with four faces. On the other hand, the second row in the first column shows that the complement of the class, namely $\{0\}^{c}:=\{1,2,3,4,5,6,7,8,9\}$, can be separated from $\{0\}$ by a convex polytope with three faces, as illustrated in Figure 6(a).
Class number
@@
Table 1: Polytope-basis covers of each class in MNIST, Fashion-MNIST, and CIFAR10 datasets. Here, $a$+$b$ denotes two polytopes with $a$ and $b$ faces. Note that each class of real-world datasets can be covered by at most two polytopes, indicating the geometric simplicity of real datasets.
For Fashion-MNIST and CIFAR10 datasets, certain classes are covered by two polytopes instead of just one. Figure 6(b) visually illustrates the classification of the class $\{0\}$ in Fashion-MNIST, accomplished through the difference of two polytopes. In the case of the CIFAR10 dataset, the number of faces in the polytopes tends to be higher compared to other datasets, consistent with the expectation that CIFAR10 exhibits a more intricate geometric structure than MNIST or Fashion-MNIST.
Furthermore, we identify the geometric complexity of each class from Table 1. In Fashion-MNIST, the complement of the class in $\{0,2,3,4,6\}$ prominently require more faces than other classes, and they all pertain to top clothes and share visual similarities (Figure 6(b)). Furthermore, it fails to find a single convex polytope cover of $\{0\}$ since the obtained polytope always contains many images in $\{2,3,4,6\}$ classes, as illustrated in Figure 6(b). In contrast, the class $\{1\}$ and $\{1\}^{c}$ are separated by the smallest number of faces, indicating that they are less entangled with other classes. This example demonstrates how neural networks can be utilized as a tool for quantifying the geometric complexity of datasets. We provide further empirical examples in Appendix B.
@@
By combining the results in Table 1 with Theorem 3.4, we can ascertain the feasible architectures of these datasets, based on their geometric characteristics. Specifically, adopting the covering polytopes with minimal number in Table 1 for each class, we deduce that
@@
are feasible architectures for these datasets. Note that this is the first result on the minimal architectures that can completely classify the given datasets, utilizing the geometric features of the datasets.
It is worth noting that our findings stem from Algorithm 1, which selectively removes and adjusts neurons within the network. Given the sparse connectivity observed in these networks (as demonstrated in the proof of Theorem 3.4 and illustrated in Figure 2(c)), we anticipate an inherent connection between our results and the Lottery Ticket Hypothesis (Frankle & Carbin, 2018; Malach et al., 2020). Notably, in Remark 4.1, we established the existence of the minimal network architectures for these datasets. Exploring how these minimal architectures could be identified and pruned from trained networks represents an intriguing avenue for future investigation.

### v2

@@ 3.1 Data Geometry-Dependent Bounds on Widths
for some constant $M$. Then $\mathcal{N}({\bm{x}})=1$ for all ${\bm{x}}\in C$, and we can prove that $d\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}0pt\rightarrow 1$ is a feasible architecture on the polytope, by adjusting the value of $M$. The detailed proof can be found in Appendix E.1. ∎
@@
Building upon Theorem 3.4, we can further explore the relationship between the topological properties of a dataset and the maximum width achievable by feasible network architectures. Specifically, we concentrate on simplicial complexes and Betti numbers, which are fundamental tools for investigating the topological structure of point cloud datasets in topological data analysis (TDA).
A simplicial $j$-complex is a specific type of simplicial complex where the highest-dimensional simplex has dimension $j$. Within a given simplicial complex $K$, a facet is a simplex with the highest dimension that is not a face (subset) of any larger simplex (Magai & Ayzenberg, 2022). With these definitions established, we proceed by proposing a feasible network architecture and deriving upper bounds on its width when one class $\mathcal{X}_{+}$ forms a simplicial complex.
@@
The proof can be found in Appendix E.3. Theorem 3.5 reveals that the width $d_{1}$ is bounded by in terms of the number of facets $k$ of the provided simplicial complex. From a geometric perspective, it is generally intuitive that a smaller number of facets suggests a simpler structure of the simplicial complex. This notion is mathematically expressed in (3.5), which suggests that the first value in (3.5) results in $d_{1}\lesssim\frac{k}{2}(d+3)$, which magnifies as $k$ increases. Similarly, when the maximal dimension $J$ is smaller than $\frac{d}{2}$ and $k$ is fixed, the summation in the second term in (3.5) reduces to $d_{1}\lesssim(d+1)\left(k\frac{J+2}{d-J}+2\right)$, which rapidly diminishes as $J$ decreases. This analysis demonstrates that a smaller dimension $J$ demands smaller widths, which aligns with the intuition that the lower-dimensional manifold could be approximated with the smaller number of neurons.
Now, we demonstrate how the result in Theorem 3.4 can be further leveraged to ascertain a neural network architecture with width bounds defined in terms of the Betti numbers. The Betti number is a key metric used in TDA to denote the number of $k$-dimensional ‘holes’ in a data distribution, which are frequently employed to study the topological characteristics of topological spaces (Naitzat et al., 2020).
Recall that Theorem 3.4 offers an upper bound on widths when $\mathcal{X}$ can be depicted as a difference between two groups of convex sets. Expanding on this, assuming the polytope-basis cover consists of prismatic polytopes [^2] For the formal definition of prismatic polytopes, see Appendix E.4, we can derive a bound for network architecture in relation to its Betti numbers. The result is concretely explained in the following theorem.
@@
Let $\mathcal{X}=\mathcal{X}_{+}\cup\mathcal{X}_{-}$ be a union of two disjoint topological spaces, where $\mathcal{X}_{-}$ can be separated from $\mathcal{X}_{+}$ by disjoint bounded prismatic polytopes having at most $0pt$ faces. Let $\beta_{k}$ be the $k$-th Betti number of the polytope-basis cover. Then, the following three-layer architecture
@@
The proof is provided in Appendix E.4. One of the important implications of Theorem 3.6 is the upper and lower bounds on network widths in terms of the Betti numbers of $\mathcal{X}$, which reveals the interplay between the topological characteristics of the dataset and network architectures. In Appendix, we also show in Proposition F.2 that topological property alone cannot determine the feasible architecture, highlighting the significance of prismatic polytopes assumption in Theorem 3.6. In other words, the result in Proposition F.2 implies that the geometrical assumptions in this theorem are indispensable to connecting topological features with bounds on the network widths.
Interestingly, the sum of Betti numbers $\sum_{k=0}^{d}\beta_{k}$ which appears in the third layer in (3.6), is often called the topological complexity of $\mathcal{X}$. This quantity is recognized as a measure of the complexity of a given topological space in some previous works (Bianchini & Scarselli, 2014; Naitzat et al., 2020), and can be bounded by Morse theory (Milnor et al., 1963) or Gromov’s Betti number Theorem (Gromov, 1981).
Furthermore, the lower bound on widths (4) shows that the sum of product of widths should be greater than the sum of Betti numbers. This finding also confirms the increased significance of widths in deeper layers as compared to earlier ones, highlighting the advantageous impact of depth in network architecture. It also verifies that the contribution of the width in deeper layers holds greater significance compared to previous layers, i.e., the positive effect of depth.
@@ 3.2 Polytope-Basis Cover Search Algorithm
Figure 4: A polytope-basis cover derived from a trained three-layer ReLU network defined in (6), obtained from Algorithm 1. The decision boundary and activation boundaries444 For the formal definition of decision boundary (DB) and activation boundaries (ABs), see Definition C.1 in Appendix C.  of the trained network are depicted in (a). Each polytope corresponding to $a_{j}=+1$ and $a_{j}=-1$ is illustrated in (b) and (c), respectively. The result constitutes a polytope-basis cover of the swiss roll dataset.
@@
$\mathcal{T}_{j}({\bm{x}})$ $:=\lambda+\sum_{k=1}^{0pt_{j}}v_{jk}\sigma({\bm{w}}_{jk}^{\top}{\bm{x}}+b_{jk}),\quad\forall v_{jk}<0,$ (5)
$\mathcal{N}({\bm{x}})$ $:=-\frac{1}{2}\lambda+\sum_{j=1}^{J}a_{j}\sigma(\mathcal{T}_{j}({\bm{x}})),\quad\forall a_{j}\in\{\pm 1\}$ (6)
for a positive constant $\lambda$. For a given dataset ${{\mathcal{D}}}=\{({\bm{x}}_{i},y_{i})\}$, suppose $\mathcal{N}$ satisfies
$\sigma(\mathcal{T}_{j}({\bm{x}}_{i}))=0\;\textrm{ or }\;\lambda,\quad\forall{\bm{x}}_{i}\in{{\mathcal{D}}},\;\forall j\in[J].$ (7)
Then, the collection of polytopes $\{C_{j}\}_{j\in[J]}$, defined by $C_{j}:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}_{j}({\bm{x}})=\lambda\}$, becomes a polytope-basis cover of ${{\mathcal{D}}}$ whose accuracy is same with $\mathcal{N}$.
Algorithm 1 Compressing algorithm
@@
The proof of this theorem can be found in Appendix E.5. The constant $\lambda$ in (5) and (7) is a positive scalar value determined from the ratio of labels in the dataset. In experiments, we practically adopted $\lambda=5$. See Appendix C.3 for further details.
This theorem establishes that if a trained three-layer network $\mathcal{N}$ defined in (6) satisfies the condition outlined in (7), then we can derive a polytope-basis cover of the training dataset ${{\mathcal{D}}}$ from $\mathcal{N}$. Below, we outline two strategies we employed to satisfy both (5) and (6) conditions.
Firstly, to meet (5), we utilize the implicit bias of gradient descent established by Du et al. (2018, Theorem 2.1), stated in Proposition F.5. Specifically, we initialize the network weights to satisfy
@@
Then, the implicit bias preserves the inequality (8), thus ensures $v_{jk}<0$ for all $j\in[J]$ and $k\in[m]$ on the gradient flow. This satisfies the first condition.
Secondly, to achieve (7), we introduce a novel approach named the “compressing algorithm.” This algorithm aims to ‘compress’ a given two-layer network $\mathcal{T}_{j}$ defined in (5) to identify a minimal convex polytope representation. The process is precisely outlined in Algorithm 1.
More precisely, the algorithm operates in two main steps: 1) identifying and removing a neuron that do not significantly influence the decision boundary, and 2) amplifying the weights to delineate the faces of the decision boundary polytope. We illustrate the functionality of the algorithm in Figure 3. In Figure 3(a), the red arrow highlights a neuron identified as non-essential for the decision boundary. This neuron is subsequently removed as shown in (b). Subsequently, by scaling the weights $(v_{k},{\bm{w}}_{k},b_{k})$ by a factor $\lambda_{scale}>1$, the decision boundary shrinks into a convex polytope with a number of faces equal to the width of $\mathcal{T}$ (depicted in Figure 3(c)).
Figure 5: Experimental verification of convergence of gradient descent. Two columns exhibit the shape of two topological spaces, which are ‘two triangles’ and ‘a hexagon with a pentagon hole.’ The second and third row show the converged networks by gradient descent under the BCE loss and the MSE loss, respectively. The first layer (blue) and second layer (red) represent the hyperplane and polytopes, respectively, described in Section 3.
Note that a single execution of Algorithm 1 may not immediately yield the network satisfying (7). However, we prove in Proposition C.2 that by iterating this algorithm a sufficient number of times, the output of the algorithm always satisfies both (5) and (7), making it suitable for the application of Theorem 3.7. Therefore, we apply the algorithm once every thousand iterations during the gradient descent optimization process, and the end of whole training. It is precisely described in Algorithm 2 in Appendix C.1. We additionally mention that Algorithm 1 is compatible with non-pretrained networks, although this flexibility may come at the cost of increased training time.
[…]
Lastly, we mention that we further propose two alternative methods for obtaining a polytope-basis cover in Appendix C. Specifically, one approach involves deriving such a cover from any trained two-layer network (Algorithm 3), while the other method entails training several neural networks in order (Algorithm 4). Further details and comparisons of these additional algorithms are provided in Appendix C.
Figure 6: Illustration of a polytope-basis cover of the real datasets. (a) The class $\{0\}$ of MNIST can be separated by a single convex polytope with four faces, while the complement class $\{0\}^{c}$ can be separated with three faces. (b) The class $\{0\}$ of Fashion-MNIST can be separated by the difference of two polytopes, one of which contains similar images. Other classes also have simple polytope-basis covers as described in Table 1.
@@ 4 Experimental Results
In Section 3, we studied the relationship between the dataset geometry and neural network architectures. In this section, we provide two empirical results: 1) gradient descent indeed converges to the networks we unveil, and 2) we can investigate the geometric features of high-dimensional real-world datasets through our proposed algorithm.
@@ 4.1 Convergence on Polytope-Basis Covers
For the first dataset $\mathcal{X}_{1}$, Theorem 3.5 suggests that $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}6\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}2\rightarrow 1$ is a feasible architecture on $\mathcal{X}_{1}$. To facilitate a clearer examination of weight vectors in each layer, we plot the activation boundaries4 in blue (the 1st hidden layer) and red (the 2nd hidden layer) colors, where the gray color denotes the decision boundary of the converged network. The weight vectors in the first layer accurately enclose the two triangles, reflecting the geometrical shape of $\mathcal{X}_{1}$. Similarly, for the second dataset $\mathcal{X}_{2}$, Theorem 3.4 suggests that $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}11\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}2\rightarrow 1$ is a feasible architecture . Specifically, the eleven neurons in the first layer correspond to the eleven hyperplanes delineating the boundaries of the outer hexagon and the inner pentagon, while two neurons in the second hidden layer align with the two polygons. These outcomes precisely correspond to our network constructions depicted in Figure 2(b, c).
We conclude this section by providing theoretical insights into the convergence behavior of gradient descent. In Appendix D, utilizing our explicit construction of neural networks, we construct an explicit path that loss strictly decreases to zero (the global minima), when the network is initialized close to the target polytope (see Theorem D.3). The specific conditions governing the initialization region are described in terms of the distribution of the dataset along the convex polytope. Although this result does not mean that gradient descent must converge to the global minimum but only guarantees the existence of such a path, it is strong evidence for the convergence to the global minima. For a thorough understanding and precise statements, we refer readers to Appendix D.
@@ 4.2 Polytope-Basis Cover for Real Datasets
Figure 7: Visualization of two polytope covers for MNIST classes $\{0\}$ and $\{8\}$. (a) The four faces of two distinct polytope covers for class $\{0\}$ in the MNIST dataset are depicted. (b) The eight faces of two distinct polytope covers for class $\{8\}$ in the MNIST dataset are depicted. The distribution of angles between the vectors in two polytope covers are plotted on the right panel.
@@
Our empirical results are presented in Table 1. Each column in the table corresponds to a class in the dataset, where each row presents the type of the class. The values in the table denotes the number of polytopes and their faces (we use notation $a$+$b$ to denote two polytopes with $a$ and $b$ faces, respectively). For instance, the value in the first row and the first column shows that the $\{0\}$ class images in MNIST dataset can be distinguished from other classes by a single convex polytope with four faces. On the other hand, the second row in the first column shows that the complement of the class, namely $\{0\}^{c}:=\{1,2,3,4,5,6,7,8,9\}$, can be separated from $\{0\}$ by a convex polytope with three faces, as illustrated in Figure 6(a).
Class
@@
Table 1: Polytope-basis covers of each class in MNIST, Fashion-MNIST, and CIFAR10 datasets. Here, $a$+$b$ denotes two polytopes with $a$ and $b$ faces. For certain classes that cannot be covered by a single convex polytope with less than 30 faces, a second polytope is additionally computed. Overall, each class of real-world datasets can be covered by at most two polytopes, indicating the geometric simplicity of real datasets.
For Fashion-MNIST and CIFAR10 datasets, certain classes that cannot be covered by a convex polytope with less than 30 faces are covered by two polytopes. Figure 6(b) visually illustrates the classification of the class $\{0\}$ in Fashion-MNIST, accomplished through the difference of two polytopes. In the case of CIFAR10 dataset, the number of faces in the polytopes tends to be higher compared to other datasets, consistent with the expectation that CIFAR10 exhibits a more intricate geometric structure than MNIST or Fashion-MNIST.
Furthermore, we identify the geometric complexity of each class from Table 1. In Fashion-MNIST, the complement of the class in $\{0,2,3,4,6\}$ prominently require more faces than other classes, and they all pertain to top clothes and share visual similarities (Figure 6(b)). Furthermore, it fails to find a single convex polytope cover of $\{0\}$ (with less than 30 faces) since the obtained polytope always contains many images in $\{2,3,4,6\}$ classes as illustrated in Figure 6(b). In contrast, the class $\{1\}$ and its complement $\{1\}^{c}$ are separated by the fewest faces, suggesting they are less entangled with other classes. This observation is consistent with the distinctive, unique shape of the “Trouser” class in Fashion-MNIST dataset. This result highlights how neural networks can be utilized as a tool for quantifying the geometric complexity of datasets. We provide additional interesting empirical examples in Appendix B.
We further investigate the uniqueness of the obtained polytope covers in MNIST dataset. When Algorithm 1 is applied to networks initialized with small norm, the obtained covers exhibit noticeable similarity. For MNIST class $\{0\}$, we compute two distinct covering polytopes $\mathcal{C}_{0}$ and $\tilde{\mathcal{C}}_{0}$ with four faces. In Figure 7(a), the weight vectors ${\bm{w}}_{i}$ of these covers are visualized, and the angles between the vectors of these two polytopes are displayed. It is easily verified that there is a clear one-to-one correspondence between vectors in the two polytopes, both visually and numerically.
For MNIST class $\{8\}$, which has a polytope cover with eight faces, a similar result is provided in Figure 7(b). Although the correspondence is slightly weaker than that of class $\{0\}$ due to the increased number of faces, most vectors still exhibit a strong one-to-one correspondence. From this, it can be seen that Algorithm 1 experimentally provides a unique polytope cover, offering further epexegetic support for the geometric simplicity of MNIST dataset.
Now, we provide feasible architectures for multi-class classifier for real datasets. By combining the results in Table 1 with Theorem 3.4, we can ascertain the feasible architectures of these datasets, based on their geometric characteristics. Note that this is the first result on the minimal network architectures that can completely classify the given datasets, utilizing the geometric features of the datasets. It is provided in the remark below.
@@
Adopting the covering polytopes with minimal number in Table 1 for each class, we deduce that
@@
are feasible architectures for these datasets. Furthermore, the second and third weight matrices in these networks are highly sparse, as demonstrated in the proof of Theorem 3.4 and illustrated in Figure 2(c).
It is worth noting that our findings stem from Algorithm 1, which selectively removes and adjusts neurons within the network. Given the sparse connectivity observed in these networks, we anticipate an inherent connection between our results and the Lottery Ticket Hypothesis (Frankle & Carbin, 2018; Malach et al., 2020). In other words, the sparse pruned neural networks suggested in Remark 4.1 can be understood as an example of the ‘winning ticket’ in LTH that is explicitly constructed.
Additionally, our results offer another perspective on understanding LTH. For instance, the assumptions in Theorem 3.7 shed light on the significance of masked and unmasked weights, and why maintaining the sign of weight values is important (Zhou et al., 2019). Specifically, the unmasked weights can be associated with the connection of faces to polytopes, and preserving the signs is crucial for maintaining the convex structure of these polytopes, as specified in (5) and (7). We hope our study contributes to future research efforts aimed at elucidating the principles underlying LTH.